luismaya:

Eliza started her savings account with $100. Each month she deposits $25 into her account. Determine the average rate of change in Eliza's account from the 2nd month to the 10th month. Show all work for full credit.

luismaya:

i need help with this answer . if some one could kindly help. thank you

surjithayer:

amount=100+25x , where x is the number of months. after 1 month or amount in second month=100+25*1=125 amount after 9 months or in 10 th month=100+25*9=100+225=325 \[average ~ change ~from~ 2nd~ \to ~10 th~month=\frac{ 325-125 }{ 9-1 }=\frac{ 200 }{8}=25\]

surjithayer:

also as the amount increases 25 $ every month we can straight way say average rate of change per month is 25